If the demand functions for the two consumers are:

For consumer 1 Qd1= 10-4P1

For consumer 2 Qd2= 4-2P2

When the two prices are 3 birr.

a. Determine the market demand using the same price for

the two functions?


1
Expert's answer
2022-01-14T13:51:44-0500

Q=Qd1+Qd2=10-P1+4-P2

P2=P1=3

Q=Qd1+Qd2=10-P1+4-P2=14-6=9